According to indexmundi.com China produces the largest amount of ammonia with 45,520 in thousand metric ton. The profitability of the ammonia production is around 90% of the total ammonia production cost which is about Php 10,000 per ton.

Refer to the balanced chemical equation: N2(g) + 3H2(g) ➡️ 2NH3(g)


How much nitrogen and hydrogen were used in the production of ammonia in China?


1
Expert's answer
2021-11-15T10:58:34-0500

The production cost of one ton of ammonia = 10000/1.9 = 5263.2 php per ton.

The production cost of the largest amount of ammonia = 5263.2×45,520,000 = 239,580,864,000 php.

Answer: The production cost of the largest amount of ammonia is 239,580,864,000 php